# P. Guiraud, M. Hernández, A. Meyroneinc and A. Nogueira

### v1

@@ 6.2.2 Image of $\phi$ and attractor of $f$ for $\rho$ rational
Now, for any interval of the form $[b,c)\subset[0,1)$ such that $\{\eta,a\}\cap(b,c)=\emptyset$, we have $f([b,c))=[f(b),f(c^{-}))$ with $f(c^{-})=f(c)$ if $c\notin\{\eta,a\}$. Also, on one hand, either $\eta\in(\gamma_{l},\sigma_{l+1})$ or $\eta=\sigma_{l+1}$ and on the other hand, either $a\in(\sigma_{s},\gamma_{s})$ or $a=\gamma_{s}$. Then, from (61) and (66), we deduce that
@@
which proves (62). Also, $[\sigma_{s},\gamma_{s})=[\sigma_{s},a)\cup[a,\gamma_{s})$ and $f([\sigma_{s},\gamma_{s}))=[\sigma_{j_{s}},f(a^{-}))\cup[f(a),\gamma_{j_{s}})\subset[\sigma_{j_{s}},\gamma_{j_{s}})$, since $f(a^{-})<f(a)$. Analogously,
@@
In the sequel of the proof $J:=[b,c)\neq\emptyset$ is one of the interval $[\sigma_{s},a),[a,\gamma_{s}),[\gamma_{l},\eta),[\eta,\sigma_{l+1})$ and $(c_{k})_{k\in\mathbb{N}}$ is the sequence defined by $c_{0}=c$ and $c_{k+1}=f(c_{k}^{-})$ for every $k\in\mathbb{N}$.
@@
$[f^{k+1}(b),c_{k+1})\subset[\sigma_{s},\gamma_{s}).$ (70)
@@
Let $\delta\in(1-\lambda-d,1)$. We first we show that $\rho_{\delta}$ is well defined. Let $\alpha\in[0,1]$ and consider the set
@@
This implies that there exits $m_{0}\in\mathbb{N}$ such that
@@ A.1 Continuity of the rotation number
Let :=$\mathcal{M}:=\{F:\mathbb{R}\to\mathbb{R}:\text{ $F$ is non-decreasing and $F(t+1)=F(t)+1$ $\forall t\in\mathbb{R}$}\}$. For any $F\in\mathcal{M}$ let $R(F)$ be defined by
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ A.2 Semi-conjugacy for irrational rotation numbers
Let $x\in\overline{\Lambda}\setminus\Lambda$. Note that $0=\{\phi(1^{-})\}$ is an accumulation point of $\Lambda$ such that $0=\phi(0)\in\Lambda$, since $\Lambda=\phi([0,1))$ by Theorem 2.9. So $x\neq 0$. As $\phi$ is increasing and right continuous, it follows that $x=\phi(y^{-})$ for some $y\in[0,1)$ with $\phi(y^{-})<\phi(y)$, that is, $y$ is a point of discontinuity of $\phi$. By definition, $f$ is discontinuous at 00 and $a$ with the circle topology, so it is continuous at $x$ unless $x=a$. If $D=\emptyset$, then $a(\delta,\rho^{+},\alpha)=a=a(\delta,\rho,\alpha)$ by Lemma 4.5 and (20). So $\phi(\alpha^{-})=a=\phi(\alpha)$ by (40) and it follows that $x\neq a$. If $D\neq\emptyset$, then $D\cap\mathbb{Z}^{+}\neq\emptyset$ and $\phi(\alpha^{-})=a(\delta,\rho^{+},\alpha)<a(\delta,\rho,\alpha)=\phi(\alpha)$ by Lemma 4.5 and (40). Since $a\neq a(\delta,\rho^{+},\alpha)$ by hypothesis, we deduce also that $x\neq a$. Therefore, $f$ is continuous at $x$.
@@
Case 2: Suppose that $D\cap\mathbb{Z}^{-}\neq\emptyset$. Then $\delta(\rho^{-},\alpha)<\delta(\rho,\alpha)$ by Lemma 4.5. So $0\leq\phi(0)$ and $\phi(1^{-})=\phi(0^{-})+1<1$ by (38), (20) and the hypothesis $\delta\neq\delta(\rho,\alpha)$. Then $[\phi(0),\phi(1^{-})]\subset[0,1)$ and the induced euclidean and circle topology in $[\phi(0),\phi(1^{-})]$ are the same. Consider the function $\varphi^{\prime}:[\phi(0),\phi(1^{-})]\to[0,1]$ defined by $\varphi^{\prime}(x)=\inf\{y\in[0,1]:\phi(y)\geq x\}$ for all $x\in[\phi(0),\phi(1^{-})]$. Note that $\varphi^{\prime}|_{[\phi(0),\phi(1^{-}))}=\tilde{\varphi}$ and $\varphi^{\prime}(\phi(1^{-}))=1$. This function $\varphi^{\prime}$ is also continuous in the euclidean topology. Let us consider the projection $\pi:\mathbb{R}\to[0,1)$ defined by $\pi(x)=\{x\}$ for all $x\in\mathbb{R}$. Then $\pi\circ\varphi^{\prime}:[\phi(0),\phi(1^{-})]\to[0,1)$ is continuous in the circle topology. In particular, $\bar{\varphi}:=\pi\circ\varphi^{\prime}|_{\overline{\Lambda}}:\overline{\Lambda}\to[0,1)$ is continuous. As $(\pi\circ\varphi^{\prime})|_{\Lambda}=\pi\circ\tilde{\varphi}=\tilde{\varphi}$, it follows that $\bar{\varphi}$ is surjective and $\bar{\varphi}\circ f|_{\Lambda}=R_{\rho}\circ\bar{\varphi}|_{\Lambda}$ by Lemma A.4.

### v3

@@ 6.2.2 Image of $\phi$ and attractor of $f$ for $\rho$ rational
Now, for any interval of the form $b,c)\subset[0,1)$ such that $\{\eta,a\}\cap(b,c)=\emptyset$, we have $f([b,c))=[f(b),f(c^{-}))$ with $f(c^{-})=f(c)$ if $c\notin\{\eta,a\}$. Also, on one hand, either $\eta\in(\gamma_{l},\sigma_{l+1})$ or $\eta=\sigma_{l+1}$ and on the other hand, either $a\in(\sigma_{s},\gamma_{s})$ or $a=\gamma_{s}$. Then, from (65) and (70), we deduce that
@@
which proves (66). Also, $\sigma_{s},\gamma_{s})=[\sigma_{s},a)\cup[a,\gamma_{s})$ and $f([\sigma_{s},\gamma_{s}))=[\sigma_{j_{s}},f(a^{-}))\cup[f(a),\gamma_{j_{s}})\subset[\sigma_{j_{s}},\gamma_{j_{s}})$, since $f(a^{-})<f(a)$. Analogously,
@@
In the sequel of the proof $J:=[b,c)\neq\emptyset$ is one of the interval $\sigma_{s},a),[a,\gamma_{s}),[\gamma_{l},\eta),[\eta,\sigma_{l+1})$ and $(c_{k})_{k\in\mathbb{N}}$ is the sequence defined by $c_{0}=c$ and $c_{k+1}=f(c_{k}^{-})$ for every $k\in\mathbb{N}$.
@@
$f^{k+1}(b),c_{k+1})\subset[\sigma_{s},\gamma_{s}).$ (74)
@@
Let $\delta\in(1-\lambda-d,1)$. We first show that $\rho_{\delta}$ is well defined. Let $\alpha\in[0,1]$ and consider the set
@@
This implies that there exists $m_{0}\in\mathbb{N}$ such that
@@ A.1 Continuity of the rotation number
Let $\mathcal{M}:=\{F:\mathbb{R}\to\mathbb{R}:\text{ $F$ is non-decreasing and $F(t+1)=F(t)+1$ $\forall t\in\mathbb{R}$}\}$. For any $F\in\mathcal{M}$ let $R(F)$ be defined by
@@
Let $(F_{\delta,a})$ be a family of functions in $\mathcal{M}$, where $(\delta,a)$ belongs to a subset $M$ of some normed space. For $(\delta_{0},a_{0})\in M$, we say that $F_{\delta,a}$ is $H$-convergent to $F_{\delta_{0},a_{0}}$ as $(\delta,a)\to(\delta_{0},a_{0})$, and we write $F_{\delta,a}\xrightarrow{H}F_{\delta_{0},a_{0}}$, if $R(F_{\delta,a})$ converges to $R(F_{\delta_{0},a_{0}})$ in the Hausdorff metric as $(\delta,a)\to(\delta_{0},a_{0})$. Then, we have the following:
@@
[29, Theorem 5.8] Suppose that $F_{\delta,a}$ is in $\mathcal{M}$ for all $(\delta,a)$ and that $F_{\delta,a}\xrightarrow{H}F_{\delta_{0},a_{0}}$ as $(\delta,a)\to(\delta_{0},a_{0})$. If $F_{\delta_{0},a_{0}}$ is continuous or $F_{\delta_{0},a_{0}}$ is strictly increasing then
@@
[29, Lemma 3.3] Let $(F_{\delta,a})$ be a family of functions in $\mathcal{M}$. Then $F_{\delta,a}\xrightarrow{H}F_{\delta_{0},a_{0}}$ as $(\delta,a)\to(\delta_{0},a_{0})$ if and only if for each point $x_{0}$ of continuity of $F_{\delta_{0},a_{0}}$ and for each $\epsilon>0$ there exists $\xi>0$ such that for every $(\delta,a)$ satisfying $\|(\delta,a)\|<\xi$ there exists $x_{\delta,a}$ such that
@@
For any $(\delta_{0},a_{0})\in M$, we have $F_{\delta,a}\xrightarrow{H}F_{\delta_{0},a_{0}}$ as $(\delta,a)\to(\delta_{0},a_{0})$ and
@@
Let $(\delta_{0},a_{0})\in M$, $x_{0}\in\mathbb{R}$ be a continuity point of $F_{\delta_{0},a_{0}}$ and $\epsilon>0$. As $F_{\delta_{0},a_{0}}$ is discontinuous at $a_{0}$, we have $\{x_{0}\}\neq a_{0}$ and $\xi:=\min\{|\{x_{0}\}-a_{0}|/2,\epsilon\}>0$. Let $(\delta,a)\in M$ be such that $\|(\delta,a)-(\delta_{0},a_{0})\|:=\max\{|\delta-\delta_{0}|,|a-a_{0}|\}<\xi$. Then $\{x_{0}\}<\min\{a,a_{0}\}$ or $\{x_{0}\}>\max\{a,a_{0}\}$, which implies $\theta_{a}(\{x_{0}\})=\theta_{a_{0}}(\{x_{0}\})$. Thus, $|F_{\delta,a}(x_{0})-F_{\delta_{0},a_{0}}(x_{0})|=|\delta-\delta_{0}|<\epsilon$. Taking $x_{\delta,a}=x_{0}$ we obtain (85) and it follows that $F_{\delta,a}\xrightarrow{H}F_{\delta_{0},a_{0}}$ as $(\delta,a)\to(\delta_{0},a_{0})$. Now, as $F_{\delta_{0},a_{0}}$ is strictly increasing, the limit (86) follows from Theorem A.1. ∎
@@ A.2 Semi-conjugacy for irrational rotation numbers
Let $x\in\overline{\Lambda}\setminus\Lambda$. Note that $0=\{\phi(1^{-})\}$ is an accumulation point of $\Lambda$ such that $0=\phi(0)\in\Lambda$, since $\Lambda=\phi([0,1))$ by Theorem 2.10. So $x\neq 0$. As $\phi$ is increasing and right continuous, it follows that $x=\phi(y^{-})$ for some $y\in[0,1)$ with $\phi(y^{-})<\phi(y)$, that is, $y$ is a point of discontinuity of $\phi$. By definition, $f$ is discontinuous at 0 and $a$ with the circle topology, so it is continuous at $x$ unless $x=a$. If $D=\emptyset$, then $a(\delta,\rho^{+},\alpha)=a=a(\delta,\rho,\alpha)$ by Lemma 4.7 and (24). So $\phi(\alpha^{-})=a=\phi(\alpha)$ by (44) and it follows that $x\neq a$. If $D\neq\emptyset$, then $D\cap\mathbb{Z}^{+}\neq\emptyset$ and $\phi(\alpha^{-})=a(\delta,\rho^{+},\alpha)<a(\delta,\rho,\alpha)=\phi(\alpha)$ by Lemma 4.7 and (44). Since $a\neq a(\delta,\rho^{+},\alpha)$ by hypothesis, we deduce also that $x\neq a$. Therefore, $f$ is continuous at $x$.
@@
Case 2: Suppose that $D\cap\mathbb{Z}^{-}\neq\emptyset$. Then $\delta(\rho^{-},\alpha)<\delta(\rho,\alpha)$ by Lemma 4.7. So $0\leq\phi(0)$ and $\phi(1^{-})=\phi(0^{-})+1<1$ by (42), (24) and the hypothesis $\delta\neq\delta(\rho,\alpha)$. Then $\phi(0),\phi(1^{-})]\subset[0,1)$ and the induced euclidean and circle topology in $\phi(0),\phi(1^{-})]$ are the same. Consider the function $\varphi^{\prime}:[\phi(0),\phi(1^{-})]\to[0,1]$ defined by $\varphi^{\prime}(x)=\inf\{y\in[0,1]:\phi(y)\geq x\}$ for all $x\in[\phi(0),\phi(1^{-})]$. Note that $\varphi^{\prime}|_{[\phi(0),\phi(1^{-}))}=\tilde{\varphi}$ and $\varphi^{\prime}(\phi(1^{-}))=1$. This function $\varphi^{\prime}$ is also continuous in the euclidean topology. Let us consider the projection $\pi:\mathbb{R}\to[0,1)$ defined by $\pi(x)=\{x\}$ for all $x\in\mathbb{R}$. Then $\pi\circ\varphi^{\prime}:[\phi(0),\phi(1^{-})]\to[0,1)$ is continuous in the circle topology. In particular, $\bar{\varphi}:=\pi\circ\varphi^{\prime}|_{\overline{\Lambda}}:\overline{\Lambda}\to[0,1)$ is continuous. As $(\pi\circ\varphi^{\prime})|_{\Lambda}=\pi\circ\tilde{\varphi}=\tilde{\varphi}$, it follows that $\bar{\varphi}$ is surjective and $\bar{\varphi}\circ f|_{\Lambda}=R_{\rho}\circ\bar{\varphi}|_{\Lambda}$ by Lemma A.4.